OpenStudy (madeleine2):

The y-intercept of the line whose equation is 2x + y = 4 is a. -2 b. 2 c. 4

OpenStudy (seratul):

Do you know how to make it into a y=mx+b form?

OpenStudy (tkhunny):

x = 0 ==> Substitute. y-intercept appears.

Directrix (directrix):

@madeleine2 Did you get an answer for this?

OpenStudy (3mar):

steps: 1. bring y in one side and collect others in the other side 2. make sure the coefficient of y is unity 3. the free term (always a number or rational, whatever) is the value at which the line intersect with y-axis. (if +ve, it intersects at \[y^+\] , if -ve, it intersects at \[y^-\]) 

Directrix (directrix):

The y-intercept is where the line crosses the y-axis. All points on the y-axis have an x-coordinate of 0. Take the equation: 2x + y = 4 Let x = 0. 2*0 + y = 4 Solve for y. That value is the y intercept. You are here now. Tell us what you get for the value of y. @madeleine2

Directrix (directrix):

Your Task: Solve for y: --> 2*0 + y = 4